Hanna:

Algebra 2 question. Only one question, please help!!!!!!!!!!!

Hanna:

1 attachment
Vocaloid:

Treat the solid black line as a regular equation first. Find the slope using the slope formula, and the y-intercept Since we have a solid line with the shading above, replace y = with y ≥

Hanna:

I have been doing this kind of work for a few days and I still don't understand how to get the equation

Karla:

You may watch this https://youtu.be/eHPTyYbNmx4 Or https://youtu.be/hh6RAEPlza4

Karla:

For finding equation of the line you have to find its gradient okay Formula for gradient is (y2-y1)/(x2-x1) Take any 2coordinates set from graph line For coordinate the first in the bracket is x-coordinate while the right one is y-coordinate (x-coordinate ,y-coordinate) Like in this question when x-coordinate (see horizontally) is 2; the y-coordinate is (see vertically where is the line ) also 2 So one coordinate of line is (2,2) See another (0,-6) So your coordinates are (X1,y1) ; (x2,y2) (2,2) ; (0,-6) So put the values in gradient y2=-6 y1=2 x2=0 x1=2 Gradient= (-6-2)/(0-2) =4 For finding equation the formula is y-y1 = m(x-x1) In this m=gradient And x and y will remain x and y Put values in equation y-2=4(x-2) Simplify y-2=4x-8 y=4x-8+2 y=4x-6 So this is the equation y=4x-6 So as vocaloid said replace y= with y ≥ Your answer for this question will be y ≥ 4x-6